# Beyond English: Evaluating Automated Measurement of Moral Foundations in Non-English Discourse with a Chinese Case Study

### v2

@@ Methods
We use Google Translate as an example of a machine translation approach due to its accessibility and consistent performance across domains. First, we machine translate benchmarking datasets except MFV to Chinese using the Google Cloud API—Basic Translation service. Then we estimate the MF values from translated documents with established English measurements, including lexicons MFD (Graham, Haidt, and Nosek, 2009), MFD2 (Frimer et al., 2019), eMFD (Hopp et al., 2021) and MoralStrength (Araque, Gatti, and Kalimeri, 2020); word embeddings with FrameAxis (Kwak et al., 2021); and specialized-fine-tuned language models MoralBert (Preniqi et al., 2024) and Mformer (Nguyen et al., 2024).
For MFD, MFD2, and eMFD, we calculate word frequencies using the eMFDscore Python package (Hopp et al., 2021). For MFD and MFD 1.0, each document’s MF value is determined by the most frequent MF class in the respective dictionary. A document is mapped to multiple classes if there is an equal number of class matches. If there are no matching words, no class is assigned to the document. eMFD, however, assigns probabilities to its vocabulary, representing their likelihood of being associated with certain MF classes. We sum the probabilities and label the document with the class that has the highest sum.
We use the moralstrength package (Araque, Gatti, and Kalimeri, 2020) for the MoralStrength dictionary. We first test its performance of the lexicon features alone with bag-of-word methods; then we train a Support Vector Machine (SVM) model and combine its lexicon features. Since English training sets contain many non-moral labels, while the benchmarking dataset contains only moral labels, we train two SVM models: one with the full training data and another with only moral-labeled training data.
For word embedding methods, we use FrameAxis Python package (Kwak et al., 2021) to compute anchor micro-frames based on different MFDs with the word2vec embedding model (Mikolov et al., 2013). For each MFD, we generate the corresponding micro moral frames from the vocabulary in its class. We then compute and aggregate word contributions to each microframe in the document, and label the document’s moral class by identifying significant microframes through comparison with a null model.
@@
We apply C-MFD2 (Cheng and Zhang, 2023) to evaluate the performance of the local language dictionary approach. We test two techniques for locally-developed MF lexicons: word counts and embeddings. For the word embedding method, we test two approaches with the fastText model (Grave et al., 2018). One involves measure simple semantic similarity. Words in C-MFD2 are grouped into five pseudo-documents based on their MF labels, each serving as anchor frames. MF values are then determined by calculating the semantic distance between the text to be classified and these anchor frames. The other uses FrameAxis to construct anchor frames. As C-MFD2 does not have the virtue/vice dimension which is essential to calculate micro-frames in FrameAxis, we automatically assign this dimension using the RoBERTa-based Chinese sentiment model c2-roberta-base-finetuned-dianping-chinese from HuggingFace [^7] Accessed from https://huggingface.co/liam168/c2-roberta-base-finetuned-dianping-chinese.
Multilingual Language Models
We select XLM-T  [^8] Accessed from https://huggingface.co/cardiffnlp/twitter-xlm-roberta-base as the base model to test transfer learning on the multilingual language model approach (Barbieri, Anke, and Camacho-Collados, 2022). Fine-tuned on 198 million multilingual tweets on the XLM-RoBERTa architecture—originally trained on 2.5 TB of Common Crawl data (Conneau et al., 2019)—this model is particularly well-suited for analyzing online content. Also, previous research indicates that it has strong performance in cross-language deductive coding tasks such as hate speech detection, compared to other multilingual language models (Röttger et al., 2022).
[…]
@@
For the LLMs approach, we select the Llama3.1-8b instruct model, an open-source multilingual LLM developed by Meta with a reasonable balance between model performance and computational cost. Compared to closed-source models like GPT-4, Llama3.1 offers greater control, flexibility, transparency, and reproducibility—all of which are important in human-value measurement tasks.
@@ Results
Multilingual Language Models
The multilingual language model XLM-T, trained on the same English annotated data, shows a moderate but reduced performance compared to the monolingual English model Mformer. In Table 4, XLM-T shows moderate performance across all benchmarking datasets ($F1_{MFV}=0.71,F1_{CCS}=0.71,F1_{CCV}=0.63$) outperforming both machine translation and local lexicon approaches. In addition, its performance is generally consistent across the five foundation values, showing a better reliability in the fine-grained MF measurements. This consistency is likely due to training five separate classifiers. In the real-world CCV dataset, the fine-tuned XLM-T model has moderate performance in four out of five foundation values with an average F1 score of $0.63$. The “authority” model has lower performance with $F1=0.46$.
@@
Table 4: Multilingual language models for moral foundation measurement. XLM-T Avg refers to the average F1 score across five foundation models.
@@
However, the amount of annotated non-English data required to reach reliable classification thresholds for MFs is much more than for hate speech detection (Röttger et al., 2022). On average, 16 batches are needed to reach the F1 score of $0.70$ and 49 batches to reach $0.80$. The “care” model, which is the best performing model of the five, requires six additional batches to reach an F1 score of $0.70$ and 27 batches to reach $0.80$. There are 10 batches of data available for the “sanctity” model, and it shows no significant improvement even once all 10 batches are used for training.
It is worth noting that limited local language annotated data may also negatively impact the model’s performance. For example, the “fairness” model’s F1 score initially drops from $0.50$ to $0.46$ when fine-tuned with local-language data. It only begins to stabilize and improve after 11 batches. In sum, while moderate performance can be achieved with local-language annotations, considerably more data is required to attain robust classification performance with the multilingual language model approach in cross-language MF measurements.
Figure 1: Accumulated fine-tuning of the multilingual language model XLM-T with local language annotated data from the CCV dataset. Each batch includes 100 items.
Large Language Models
Table 5 shows the results of using Llama3.1-8b. With only prompt engineering, Llama3.1-8b already performs better than the machine translation and local language lexicon approaches across all benchmarking datasets. After fine-tuning with annotated English-language data, it immediately achieves strong performance on the MFV ($F1=0.81$) and CCS ($F1=0.82$) datasets, and moderate performance on the CCV ($F1=0.65$) dataset. This performance exceeds that of XLM-T with the same English-language training data ($F1=0.63$).
Additionally, we notice that English prompts generally outperform non-English prompts on Llama3.1-8b, even when analyzing non-English documents. This difference is more pronounced in the base instruct model but becomes less significant with progressive fine-tuning using annotated data. After fine-tuning with English data, the difference in the F1 scores for English and Chinese prompts on the CCV dataset decreases from 0.11 to 0.08, further reducing to 0.03 with data augmentation and 0.02 when fine-tuned with both English and augmented data.
@@
We find traditional approaches, such as machine translation and local language lexicons, may not be the most effective solutions for cross-language moral foundation measurements. Instead, fine-tuning multilingual language models and leveraging LLMs through transfer learning emerge as promising alternatives. Notably, LLMs demonstrate strong performance and data efficiency, making them particularly well-suited for addressing the challenges of cross-language MF analysis. In this section, we extend the discussion to the general human value elements measurements in computational social science research.
Above all, the machine translation approach has proven unreliable in cross-language MF measurement. Even with state-of-the-art English tools, machine translation remains suboptimal in estimating MFs from real-world non-English corpora and, in some cases, performs worse than chance. This issue is particularly concerning for measuring culturally significant human values. For example, within MFT, previous studies highlight East–West cultural differences in emphasis on various foundation values (Graham et al., 2011). Participants from Eastern regions (South Asia, East Asia, and Southeast Asia) generally place a stronger emphasis on “loyalty” compared to their Western counterparts (United States, Canada, United Kingdom, and Western Europe) (Graham et al., 2013). In our test, however, Mformer’s “loyalty” classification on translated data ($F1=0.21$) falls below random chance ($F1=0.23$), suggesting substantial loss of this cultural nuance. Our results caution against over reliance on machine translation in cross-language MF applications and raise broader methodology concerns for its reliability in measuring other cross-language human values.
@@
Researchers should also be cautious with local language lexicons—they perform worse than the machine translation for cross-language MF measurement in some cases. Given the inherent limitations of lexicon-based methods and the extensive resources required to develop them, they are inefficient for cross-language MF measurement, particularly at the document level where semantic complexity is higher. However, culturally specific moral lexicons may still provide valuable insights at the vocabulary level for cross-cultural research, though broader applications should be approached cautiously and with extra validations.
@@
Our evaluation with multilingual language models and LLMs also reveals several key implications. Above all, fine-tuning LLMs is more data-efficient than fine-tuning multilingual language models for cross-language MF measurement. In the CCV benchmarking dataset, XLM-T requires on average more than 2,000 local-language annotated records (20 batches) to achieve strong performance (i.e., $F1=0.75$). In contrast, Llama3.1-8b can reach comparable performance with only English data machine-translated to Chinese and thus no local language labeling.
In addition, data augmentation shows strong potential in enhancing LLMs’ performance on culturally distinct MF values. For example, Llama3.1-8b notably improves on “loyalty” when fine-tuned with English data machine-translated to Chinese ($F1=0.69$), compared to its performance with the original English training data ($F1=0.57$). Such data augmentation may serve as a practical and cost-effective way to improve model performance in cross-language measurements, especially in cases where mass-labeling of local-language data is infeasible due to the resource-intensive process of creating high-quality language-specific labeled datasets.
Our findings also suggest that larger or more advanced LLMs such as Llama3.1-70b may perform better for measuring MFs in non-English corpora. That is, updating the base LLM to a more power model with enhanced multilingual capabilities could further improve the effectiveness of the LLM approach for cross-language MF measurement.
@@
Nevertheless, several limitations of using LLMs for cross-language MF measurement should be acknowledged. Firstly, their performance on fine-grained MF values can vary. For example, although the measurement on some culturally sensitive values like “loyalty” is relatively reliable, others like “authority” and “sanctity” still miss significant cultural nuances. Careful prompt crafting and fine-tuning LLMs for binary instead of multi-label classification task may help mitigate this limitation.
Secondly, the model coverage remains a challenge. Our tested four Llama models in Table 6 on average successfully detect 72.50% of the MF values in the CCV dataset, with coverage varying with the prompting language and fine-tuning data. This issue can be partially attributed to the model safety policies as Llama refuses to respond to certain prompts. Some documents in the benchmarking dataset contain morally controversial content that may prevent Llama from generating outputs—a limitation that is less prominent in multilingual language models and other approaches.
@@
Thirdly, fine-tuning LLMs with limited local-language annotated data might backfire. While training with local-language annotated data is commonly used to incorporate cultural specificity and enhance transfer learning, this approach appears less effective for LLMs when data volume is small. We fine-tuned Llama3.1-8b using 20 batches of CCV data with the same strategy used in XLM-T to test data efficiency. Each batch containing 50 MF labels evenly distributed across five classes. As shown in Figure 2, The initial drop in model performance struggles to recover within the available data for care, fairness and sanctity classifiers. Researchers with limited local-language annotated data should be cautious about potential performance degradation when fine-tuning LLMs for cross-language MF measurements.
@@
Figure 2: Accumulated fine-tuning multilingual language model Llama3.1-8b with local language annotated data from the Chinese CV dataset. Each batch includes 50 items evenly distributed across five classes.
In addition, it is important to acknowledge potential limitations in our mapping scheme. The CCV dataset was originally labeled based on Chinese core values, and the mapping process to moral foundation framing may have led to some cultural information loss.
@@
Finally, we would like to highlight the essential role of human validation in measuring cross-language MFs with LLMs. Given the limitations discussed, human validation is strongly recommended to ensure the reliability of LLM-based cross-language MF measurements. LLMs can effectively assist in this process by providing rationales alongside their classification outputs, supporting human evaluators and facilitating a more efficient assessment of cross-language MF measurements.
In summary, our findings suggest LLMs trained on English, combined with prompt-engineering and data augmentation can achieve robust performance with strong reliability and data efficiency in cross-language MF measurements. Local-language annotations should focus on the culturally distinct moral values to further enhance the performance.

### v3

@@ Methods
We use Google Translate as an example of a machine translation approach due to its accessibility and consistent performance across domains. First, we machine translate benchmarking datasets except MFV to Chinese using the Google Cloud API—Basic Translation service. Then we estimate the MF values from translated documents with established English measurements, including lexicons MFD (Graham, Haidt, and Nosek 2009), MFD2 (Frimer et al. 2019), eMFD (Hopp et al. 2021) and MoralStrength (Araque, Gatti, and Kalimeri 2020); word embeddings with FrameAxis (Kwak et al. 2021); and specialized-fine-tuned language models MoralBert (Preniqi et al. 2024) and Mformer (Nguyen et al. 2024).
For MFD, MFD2, and eMFD, we calculate word frequencies using the eMFDscore Python package (Hopp et al. 2021). For MFD and MFD 1.0, each document’s MF value is determined by the most frequent MF class in the respective dictionary. A document is mapped to multiple classes if there is an equal number of class matches. If there are no matching words, no class is assigned to the document. eMFD, however, assigns probabilities to its vocabulary, representing their likelihood of being associated with certain MF classes. We sum the probabilities and label the document with the class that has the highest sum.
We use the moralstrength package (Araque, Gatti, and Kalimeri 2020) for the MoralStrength dictionary. We first test its performance of the lexicon features alone with bag-of-word methods; then we train a Support Vector Machine (SVM) model and combine its lexicon features. Since English training sets contain many non-moral labels, while the benchmarking dataset contains only moral labels, we train two SVM models: one with the full training data and another with only moral-labeled training data.
For word embedding methods, we use the FrameAxis Python package (Kwak et al. 2021) to compute anchor micro-frames based on different MFDs with the word2vec embedding model (Mikolov et al. 2013). For each MFD, we generate the corresponding micro moral frames from the vocabulary in its class. We then compute and aggregate word contributions to each microframe in the document, and label the document’s moral class by identifying significant microframes through comparison with a null model.
@@
We apply C-MFD2 (Cheng and Zhang 2023) to evaluate the performance of the local language dictionary approach. We test two techniques for locally-developed MF lexicons: word counts and embeddings. For the word embedding method, we test two approaches with the fastText model (Grave et al. 2018). One involves measure simple semantic similarity. Words in C-MFD2 are grouped into five pseudo-documents based on their MF labels, each serving as anchor frames. MF values are then determined by calculating the semantic distance between the text to be classified and these anchor frames. The other uses FrameAxis to construct anchor frames. As C-MFD2 does not have the virtue/vice dimension, which is essential to calculate micro-frames in FrameAxis, we automatically assign this dimension using the RoBERTa-based Chinese sentiment model c2-roberta-base-finetuned-dianping-chinese from HuggingFace. [^8] Accessed from https://huggingface.co/liam168/c2-roberta-base-finetuned-dianping-chinese
Multilingual Encoder-only Models
We select XLM-T [^9] Accessed from https://huggingface.co/cardiffnlp/twitter-xlm-roberta-base as the base model to test transfer learning on the multilingual encoder-only model approach (Barbieri, Anke, and Camacho-Collados 2022). Fine-tuned on 198 million multilingual tweets on the XLM-RoBERTa architecture—originally trained on 2.5 TB of Common Crawl data (Conneau et al. 2019)—this model is particularly well-suited for analyzing online content. Also, previous research indicates that it has strong performance in cross-language deductive coding tasks such as hate speech detection, compared to other encoder-only models (Röttger et al. 2022).
We follow Nguyen et al. (2024)’s experience on fine-tuning Mformer with some tweaks. First, we replace the base architecture from RoBERTa-base (Liu et al. 2019) with twitter-xlm-roberta-base and tokenization is handled by the model’s built-in tokenizer (Barbieri, Anke, and Camacho-Collados 2022), with token sequences truncated to a maximum of 512. Second, we set the learning rate, epochs, and batch size to $52e-$ 5, 3, and 16 respectively following Röttger et al. (2022). Third, we opt to fine-tune five binary classifiers rather than a single multi-label classifier because binary models generally outperform multi-label models in English MF measurements (Nguyen et al. 2024). [^10] We note that in hate speech detection, Röttger et al. (2022) found no significant performance difference between binary and multi-label models when using XLM-T. Given that multi-label models require less storage and training resources, they are a viable alternative for future applications. Fourth, we adopt a conservative under-sampling strategy in the English annotated training dataset to address class imbalance, establishing a baseline for future improvements.
@@
For the decoder-only LLMs approach, we select the Llama3.1-8b instruct model, an open-source LLM developed by Meta with a reasonable balance between model performance and computational cost. Compared to closed-source models like GPT-4, Llama3.1 offers greater control, flexibility, transparency, and reproducibility—all of which are important in human-value measurement tasks.
@@
Qualitative Analysis
We conducted qualitative analysis to gain in-depth understanding of the cultural loss in moral foundation measurements across different approaches. By randomly sampling 100 mislabeled records, we compared predicted labels to ground truth (i.e., human annotated labels), focusing on MFs that are known with cultural distinctions between English and Chinese (i.e., authority, loyalty, and sanctity).
@@ Results
Multilingual Encoder-only Models
The multilingual encoder-only model XLM-T, trained on the same English annotated data, shows a moderate but reduced performance compared to the monolingual English model Mformer. In Table 4, XLM-T shows moderate performance across all benchmarking datasets ($F1_{MFV}=0.71,F1_{CCS}=0.71,F1_{CCV}=0.63$) outperforming both machine translation and local lexicon approaches. In addition, its performance is generally consistent across the five foundation values, showing a better reliability in the fine-grained MF measurements. This consistency is likely due to training five separate classifiers. In the real-world CCV dataset, the fine-tuned XLM-T model has moderate performance in four out of five foundation values with an average F1 score of $0.63$. The “authority” model has lower performance with $F1=0.46$.
@@
Table 4: Multilingual encoder-only models for moral foundation measurement. XLM-T Avg refers to the average F1 score across five foundation models.
@@
However, the amount of annotated non-English data required to reach reliable classification thresholds for MFs is much more than for hate speech detection (Röttger et al. 2022). On average, 16 batches are needed to reach the F1 score of $0.70$ and 49 batches to reach $0.80$. The “care” model, which is the best performing model of the five, requires six additional batches to reach an F1 score of $0.70$ and 27 batches to reach $0.80$. There are 10 batches of data available for the “sanctity” model, and it shows no significant improvement even once all 10 batches are used for training.
It is worth noting that limited local language annotated data may also negatively impact the model’s performance. For example, the “fairness” model’s F1 score initially drops from $0.50$ to $0.46$ when fine-tuned with local-language data. It only begins to stabilize and improve after 11 batches. In summary, while moderate performance can be achieved with local-language annotations, considerably more data is required to attain robust classification performance with the multilingual encoder-only model approach in cross-language MF measurements.
Figure 1: Accumulated fine-tuning of the XLM-T with local language annotated data from the CCV dataset. Each batch includes 100 items.
@@
Large Language Models
Table 5 shows the results of the LLMs approach. With only prompt engineering, Llama3.1-8b already performs better than the machine translation and local language lexicon approaches across all benchmarking datasets. After fine-tuning with annotated English-language data, it immediately achieves strong performance on the MFV ($F1=0.81$) and CCS ($F1=0.82$) datasets, and moderate performance on the CCV ($F1=0.65$) dataset. This performance exceeds that of XLM-T with the same English-language training data ($F1=0.63$).
Additionally, we notice that English prompts generally outperform non-English prompts on Llama3.1-8b, even when analyzing non-English documents. This difference is more pronounced in the base instruct model but becomes less significant with progressive fine-tuning using annotated data. After fine-tuning with English data, the difference in the F1 scores for English and Chinese prompts on the CCV dataset decreases from 0.11 to 0.08, further reducing to 0.03 with data augmentation and 0.02 when fine-tuned with both English and translated Chinese data.
Qualitative Analysis of Cultural Nuances
We selected MFormer and fine-tuned Llama3.1-8b-instruct with English and translated Chinese data as models to analyze cultural loss in translation and non-translation approaches. We identified three common types of information loss during the translation process (See Table 4 in the Appendix). First, idioms and slang are often mistranslated, as in Example 1, where the Chinese phrase meaning “faked or staged accident for compensation” is simply translated to its superficial meaning. Second, contextual meaning. As illustrated in Example 2, where the English translation “look towards” misses the cultural connotation of “favoring someone,” resulting in different expressed moral foundations. Third, political euphemisms lose their cultural context, as in Example 3, where the English translation fails to convey that the text refers to civil servant behavior. These information loss in the translation process often result in mislabeling in MF measurement.
We further analyzed the non-translational approach to understand how LLMs process cultural nuances. Compared to the translational approach, LLMs better recognize slang and idioms (Examples 2 and 13), but demonstrate more subtle cultural misalignment. Our analysis revealed several distinct patterns. Most notably, LLMs failed to distinguish certain Confucian morality concepts such as filial piety, which encompass family responsibilities to respect and care for elders (Examples 6, 8, 9, and 10). Native Chinese speakers tend to labeled these instances as “authority,” whereas language models often only categorized them as “care.” Additionally, we observed misalignment regarding government and local political entities. Human annotators tended to label cases involving government (Examples 3, 4) and officials (Example 11) as “loyalty,” while LLMs classified them as “authority” or “fairness.” Finally, examining misalignment in “sanctity” revealed that Chinese annotators frequently associated sanctity with corrupt politicians or role model civil servants (Examples 11 and 12), a distinction rarely captured by LLMs.
@@
We find traditional approaches, such as machine translation and local language lexicons, may not be the most effective solutions for cross-language moral foundation measurements. Instead, fine-tuning multilingual encoder-only models and leveraging LLMs through transfer learning emerge as promising alternatives. Notably, LLMs demonstrate strong performance and data efficiency, making them particularly well-suited for addressing the challenges of cross-language MF analysis.
Above all, simple machine translation approach has proven suboptimal in cross-language MF measurement. This issue is particularly concerning for measuring culturally significant values. Machine translation often introduces bias in the rendering of slangs, contextual meanings, and political euphemisms into local languages. Moreover, when relying on pretrained English classifiers such as MFormer—which is predominantly trained on English-annotated data—this cultural information loss may be further amplified. For instance, MFormer’s classification performance on the “loyalty” foundation in translated data ($F1=0.21$) falls below random chance ($F1=0.23$), indicating a substantial loss of cultural nuance. Nevertheless, it is important to note that machine translation models are rapidly evolving. Current deficiencies—such as loss of contextual nuance or culturally embedded moral cues—limit their effectiveness in this domain. Advances in context-aware and domain-adapted translation models (Jin et al. 2023; Saunders 2022), particularly those fine-tuned on moral discourse, may mitigate some of these issues, which deserves future exploration.
@@
Researchers should also be cautious with local language lexicons—they perform worse than the machine translation for cross-language MF measurement in some cases. Given the inherent limitations of lexicon-based methods and the extensive resources required to develop them, they are inefficient for cross-language MF measurement, particularly at the document level where semantic complexity is higher. However, culturally specific moral lexicons may still provide valuable insights at the vocabulary level for cross-cultural research, though broader applications should be approached with careful validations.
@@
Fine-tuning LLMs is more data-efficient than fine-tuning multilingual encoder-only models for cross-language MF measurement. In the CCV benchmarking dataset, XLM-T requires on average more than 2,000 local-language annotated records (20 batches) to achieve strong performance ($F1=0.75$). In contrast, Llama3.1-8b can reach comparable performance with only English data machine-translated to Chinese and thus no local language labeling. Our findings also suggest that larger LLMs such as Llama3.1-70b may perform better for measuring MFs in non-English corpora. That is, updating the base LLM to a more power model with enhanced multilingual capabilities could further improve the performance of the LLM approach.
@@
Nevertheless, several limitations of using LLMs for cross-language MF measurement should be acknowledged. Firstly, the performance on fine-grained MF values can vary. Although the measurement on some culturally sensitive values like “loyalty” is relatively reliable, others like “authority” and “sanctity” still miss significant cultural nuances.
@@
Second, fine-tuning LLMs might backfire. Fine-tuning with limited local-language data may degrades model performance. While training with local language annotated data is commonly used to incorporate cultural specificity and enhance transfer learning, this approach appears less effective for LLMs in MF measurements task when data volume is small. We fine-tuned Llama3.1-8b using 20 batches of CCV data with the same strategy used in XLM-T. Each batch containing 50 MF labels evenly distributed across five classes. As shown in Figure 2, the initial drop in model performance struggles to recover within the available data for ”care”, ”fairness”, and ”sanctity” values.
In addition, fine-tuning with English-annotated data may introduce extra cultural bias to LLMs. We replicated the same finetuning strategy to an Italian dataset “moralConvITA” (Stranisci et al. 2021), and observed similar significant performance drops in culturally-nuanced values like ”loyalty” and ”sanctity” (see Appendix). This signals the risk of cultural bias in English-centric LLMs. Finetuning such LLMs with English annotation data may exacerbate their bias against non-English languages, resulting in more cultural information loss in the measurement task, which echoes concerns about LLMs being shaped by values associated with WEIRD populations (Atari et al. 2023b). These findings highlight the need for culturally-sensitive fine-tuning approaches that preserve cross-cultural moral nuances.
@@
Figure 2: Accumulated fine-tuning Llama3.1-8b with local language annotated data from the Chinese CV dataset. Each batch includes 50 items evenly distributed across five classes.
Third, we note that our experiments were conducted in a single non-English language, which may limit the generalizability of our findings. Nevertheless, we hope this work can serve as a foundation for future research that extends to a broader range of non-English languages on MF measurement, particularly low-resource ones, where such methodological advancements would be especially valuable.
@@
Finally, we highlight the essential role of human validation in measuring cross-language MFs with LLMs. Given the limitations discussed, human validation is strongly recommended to ensure the reliability of LLM-based cross-language MF measurements. LLMs can effectively assist in this process by providing rationales alongside their classification outputs, supporting human evaluators and facilitating a more efficient result assessment.